A river barge, whose cross section is approximately rectangular, carries a load of grain.  The barge is 28 ft wide and 90 ft long.  When unloaded its draft (depth of submergence) is 5 ft, and with the load of grain is 7 ft.  Determine:  (a) the unloaded weight of the barge, and (b) the weight of the grain.
Answer #1
Concepts and Reason

A river barge is a flat-bottomed boat, built mainly for river and canal transport of heavy goods. Due to the load carried by it, it sinks in water by a certain amount. The total weight of the river barge (self-weight plus load carried with it) is equal to the weight of water displaced by it.

By knowing the volume of the barge and by marking the depth of submergence, it is possible to know the total load carried by it.

Fundamentals

The weight of the barge without the grain can be calculated as,

Wb=γVb{W_b} = \gamma {V_b}

Here, the weight density is γ\gamma and the submerged volume of the barge without the grain is Vb{V_b}.

The submerged volume of the barge without the grain is,

Vb=Ah{V_b} = Ah

Here, the area of the barge is A and the depth of the barge without the grain is h.

The weight of the barge with the grain can be calculated as,

Wbg=γVbg{W_{bg}} = \gamma {V_{bg}}

Here, the submerged volume of the barge with grain is Vbg{V_{bg}}.

The submerged volume of the barge with the grain is,

Vbg=Ahg{V_{bg}} = A{h_g}

Here, the depth of the barge with the grain is hg{h_g}.

(a)

Calculate the submerged volume of the barge without the grain.

Vb=Ah{V_b} = Ah

Here, the area of the barge is A and the depth of the barge without the grain is h.

Substitute (28×90)ft2\left( {28 \times 90} \right){\rm{ f}}{{\rm{t}}^2} for A and 5 ft for h.

Vb=(28×90)×5=12600ft3\begin{array}{c}\\{V_b} = \left( {28 \times 90} \right) \times 5\\\\ = 12600{\rm{ f}}{{\rm{t}}^3}\\\end{array}

Calculate the weight of the barge without the grain.

Wb=γVb{W_b} = \gamma {V_b}

Here, the weight density is γ\gamma .

Substitute 62.4lb/ft362.4{\rm{ lb/f}}{{\rm{t}}^3} for γ\gamma and 12600ft312600{\rm{ f}}{{\rm{t}}^3} for Vb{V_b}.

Wb=62.4×12600=786240lb\begin{array}{c}\\{W_b} = 62.4 \times 12600\\\\ = 786240{\rm{ lb}}\\\end{array}

(b)

Calculate the submerged volume of the barge with the grain.

Vbg=Ahg{V_{bg}} = A{h_g}

Here, the depth of the barge with the grain is hg{h_g}.

Substitute (28×90)ft2\left( {28 \times 90} \right){\rm{ f}}{{\rm{t}}^2} for A and 7 ft for h.

Vbg=(28×90)×7=17640ft3\begin{array}{c}\\{V_{bg}} = \left( {28 \times 90} \right) \times 7\\\\ = 17640{\rm{ f}}{{\rm{t}}^3}\\\end{array}

Calculate the weight of the barge with the grain.

Wbg=γVbg{W_{bg}} = \gamma {V_{bg}}

Here, the submerged volume of the barge with grain is Vbg{V_{bg}}.

Substitute 62.4lb/ft362.4{\rm{ lb/f}}{{\rm{t}}^3} for γ\gamma and 17640ft317640{\rm{ f}}{{\rm{t}}^3} for Vbg{V_{bg}}.

Wb=62.4×17640=1100736lb\begin{array}{c}\\{W_b} = 62.4 \times 17640\\\\ = 1100736{\rm{ lb}}\\\end{array}

Calculate the weight of the grain.

Wg=WbgWb{W_g} = {W_{bg}} - {W_b}

Substitute 786240 lb for Wb{W_b} and 1100736 lb for Wbg{W_{bg}}.

Wg=1100736786240=314496lb\begin{array}{c}\\{W_g} = 1100736 - 786240\\\\ = 314496{\rm{ lb}}\\\end{array}

Ans: Part a

The unloaded weight of the barge is 786240 lb.

Part b

The weight of the grain is 314496 lb.